# Splitting of Girvin-MacDonald-Platzman density wave and the nature of chiral gravitons in fractional quantum Hall effect

### v1

@@ Abstract
Introduction. Perhaps the most dramatic manifestation of strong correlations is “fractionalization of the electron,” which refers to the notion that the emergent elementary particles (that is, the weakly interacting degrees of freedom) carry quantum numbers of a split electron [1]. For example, when an electron is added to a Luttinger liquid, its spin and charge separate [2], and when a fully spin-polarized electron is added to an incompressible FQH state, it disintegrates into fractionally-charged excited CFs [3, 4]. The emergent particles have their own independent dynamics and provide a natural explanation of the otherwise puzzling phenomenology.
This Letter concerns the Girvin-MacDonald-Platzman (GMP) density wave mode of the FQH states [5, 6]. Since the work of GMP numerous experimental [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 16] and theoretical [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36] studies have addressed the neutral excitations of the FQH states. Very recent inelastic light scattering experiments by Liang et al. [37] with circularly polarized photons have measured the chirality of the long wavelength neutral excitation, which has been called “graviton” because of its spin-2 quadrupolar character in a geometric interpretation, and found it to be consistent with theoretical predictions [25, 31, 38].
@@
$S(\mbox{\boldmath$q$},E)=\sum_{\alpha}|\langle\Psi_{\nu}^{\alpha}|\Psi^{\rm GMP}_{\mbox{\boldmath$q$}}\rangle|^{2}\delta(E_{\alpha}-E_{0}-E),$ (1)
@@
In general, $E^{\rm SMA}_{\mbox{\boldmath$q$}}$ is the “average energy” of the excitations coupled to the ground state by $\bar{\rho}_{{}_{\mbox{\boldmath$q$}}}$. $\Psi^{\rm GMP}_{\mbox{\boldmath$q$}}$ can be evaluated explicitly for small systems for which the Fock space representation of the exact or the trial ground state is available but not for larger systems. GMP related the energy of this mode to the static structure factor $S(\mbox{\boldmath$q$})$, which is the Fourier transform of the pair correlation function $g(\mbox{\boldmath$r$})$, which can be evaluated for a trial wave function for fairly large systems [39, 40, 41, 42], allowing for a determination of the energy of the GMP mode [5, 6, 43, 22].
@@
We employ two methods for treating the LLL projection. In “direct projection” one expresses the wave function in the Fock basis and retains only the part that resides in the LLL [18]. This method can be implemented only for small systems, as the dimension of the Fock space grows exponentially with $N$ (see Supplemental Material (SM) [47]). The Jain-Kamilla (JK) method [48, 49, 50, 51, 52] allows treatment of much larger systems. The two methods generally produce wave functions that are extremely close but not identical.
Figure 1: (a) GMP density wave mode, created by applying projected electron density operator on the ground state. (b) Jain CF ground state at $\nu{=}2/5$. (c)-(g) Examples of single CF excitons. (h),(i) Schematic depiction of splitting of the GMP density wave mode (dashed black curve) into a ladder of CF excitons (blue dotted) which can be obtained by diagonalizing the Coulomb interaction in the CFE basis. The actual lowest neutral mode is well described as the primary CFE shown in (c). For $\nu{=}n/(2n{\pm}1)$ the GMP graviton is essentially identical to the primary CF graviton (h) while for $\nu{=}n/(4n{\pm}1)$ with $n{>}1$ the GMP graviton splits into two CF gravitons, which can be also be nicely understood as the primary CF and parton gravitons (i). (j) Parton construction of $\nu{=}2/9$. Each electron is divided into one parton of charge $e/9$ filling two LLs and four partons of charge $92e/$ 9, each filling a single LL. (k) CF exciton. (l) Parton exciton.
Spherical geometry. We use the spherical geometry [53] in which $N$ electrons move on the surface of a sphere in the presence of a radial magnetic field generated by a magnetic monopole of strength $2Qhc/e$, where $2Q$ is an integer. The LL orbitals are given by the monopole harmonics $Y_{Q,l,l_{z}}$ [54, 55], where $l{=}|Q|,|Q|{+}1,{\cdots}$ and $l_{z}{=}{-}l,{-}l{+}1,{\cdots},l$. The $n$th LL corresponds to $l{=}|Q|{+}n$, with $n{=}0,1,2{\cdots}$. The eigenstates are labeled by the total orbital angular momentum $L$ and its $z$-component $L_{z}$.
@@
In all cases studied, the GMP mode and CFEs constructed with direct or JK projection vanish identically for $L{=}1$. For higher $L$, we have studied the following ($\nu$, $N$) systems with direct projection: ($1/3$,6) for $L{\leq}7$; ($1/3$,8) for $L{\leq}7$; ($2/5$,6) for $L{\leq}4$; ($2/5$,8) for $L{\leq}5$; ($2/3$,6) for $L{\leq}4$; ($2/3$,8) for $L{\leq}4$; ($2/7$,6) for $L{\leq}4$; ($2/7$,8) for $L{\leq}4$; ($2/9$,6) for $L{\leq}4$; bosonic ($2/3$,6) for $L{\leq}4$; bosonic ($2/3$,8) for $L{\leq}4$. We find ${\cal O}{=}1$ to machine precision for all of these. This is nontrivial, because, as shown in Table 1, the number of CFEs at a given $L$ is $N$-independent (provided $N$ is sufficiently large) and much smaller than the number of basis functions, which ranges to several thousand for the systems studied and increases exponentially with $N$. This gives us confidence that the GMP mode is generally contained entirely within the (direct-projected) CFE subspace, although we have not yet succeeded in devising a proof.
@@
The values of ${\cal O}$ obtained for the JK-projected CFEs are given in Table 1 for $\nu{=}n/(2n{+}1)$. In all cases studied, we find that ${\cal O}{>}0.997$, demonstrating that Eq. (4) is practically exact for the JK projection. The JK-projected CFEs for $n/(4n{\pm}1)$ Jain states are considered below.
The relation in Eq. (4) between the GMP mode and single CFEs is unexpected as these are motivated by qualitatively different physical descriptions. The GMP ansatz creates a single electron-hole pair excitation, while a CFE involves the excitation consisting of a single CF particle and a single CF hole, each of which, in isolation, has a fractional charge of magnitude $e/(2n{\pm}1)$ [3]. Indeed, the dispersion of the GMP mode is rather different from that of the primary CFE [45, 46, 22]. Moreover, because an electron added to a FQH state is described as a complex bound state of $12n{\pm}$ 1 excited CFs [59, 60, 61, 62], one would a priori expect an electron-hole pair of the GMP mode to involve $12n{\pm}$ 1 CFEs.
CFE ansatz for $S(\mbox{\boldmath$q$},E)$. Substituting Eq. (4) into Eq. (1) suggests that the SMA of Eq. (2) can be replaced by a “CFE Ansatz” for the DSF:
@@
Figure 2: The energies of the GMP mode (green diamonds), the primary CF exciton (blue dots) for $\nu{=}1/3,~{}2/5,~{}3/7,~{}1/5,~{}2/9$ and $2/7$, and the primary parton exciton (black pentagrams) for $\nu=2/7,2/9$. Red dashes show the exact Coulomb spectrum. The spherical geometry is used, and the GMP energy is computed using the Laughlin (Jain) ground state for $1/3,~{}1/5$ ($2/5,~{}3/7,~{}2/9,~{}2/7$).
Figure 3: Plot of $(\rho(r)-\rho_{0})/\rho_{0}$ for the CF graviton with $L{=}2{=}L_{z}$ at $\nu{=}1/3$, $2/5$ and $3/7$, where $\rho(r)$ and $\rho_{0}$ are the densities of the graviton and the ground state, and $r$ is the distance from the center.
@@
For the Jain states at $\nu{=}n/(4n{\pm}1)$ with $n{>}1$ the energy of the GMP graviton differs significantly from that of the primary CF graviton, as seen in Fig. 2. Fig. 4 shows the spectral weight obtained by diagonalizing the Coulomb interaction within the direct-projected CF graviton space (which has three linearly independent modes). It is in excellent agreement with that obtained by ED, supporting the validity of Eq. (5). The presence of two prominent peaks indicates that the GMP graviton splits into two CF gravitons.
Remarkably, Nguyen and Son [67] predicted an additional high-energy graviton at $n/(2pn{\pm}1)$ for $22p{\geq}4,n{\geq}$ 2, motivated by the observation that a Dirac CF field theoretical description of $\nu{=}n/(4n{\pm}1)$ states by Goldman and Fradkin [68] and Wang [69] does not satisfy a bound derived by Haldane [70] which says that the coefficient of the $q^{4}$ term in the projected static structure factor is lower-bounded by $(\mathcal{S}{-}1)/8$, where $\mathcal{S}$ is the Wen-Zee shift [71]. Balram et. al. [34] proposed to identify the second, higher energy graviton as a “parton graviton.” In the parton construction of the FQH states  [72] one divides each electron into an odd number of unphysical fractionally charged fermions called partons [see Fig. 1(j)], places each species of partons in an integer quantum Hall state, and then combines the partons back into physical electrons. In the Jain $n/(2pn{\pm}1)$ states, one species of partons forms the $\nu{=}{\pm}n$ state and remaining $2p$ partons the $\nu{=}1$ state. Balram et. al. [34] defined the parton exciton as the exciton in $\Phi_{1}$ i.e.,
@@
Surprisingly, Eq. (4) is not a good approximation for the JK projected CFEs for the Jain $\nu{=}n/(4n{\pm}1)$ states with $n{>}1$. For example, for $\nu{=}2/7$ ($\nu{=}2/9$) we find ${\cal O}{=}0.5076$ (${\cal O}{=}0.9191$) for $N{=}8$ at $L{=}2$. [To project, we write $\Psi_{n/(4n{\pm}1)}{=}\Phi_{1}^{2}\Psi_{n/(2n{\pm}1)}$; 2/7 has a single CF graviton because 2/3 does (the hole conjugate of the single graviton at 1/3), and 2/9 inherits 2 CF gravitons from 2/5.] It turns out that the GMP graviton lies, to an excellent approximation, within the 2-dimensional subspace defined by the primary CF and parton gravitons. For the $N{=}8$ particle $\nu{=}2/7$ ($\nu{=}2/9$) state, the GMP graviton has ${\cal O}{=}0.9995$ (${\cal O}{=}0.9986$) with this subspace. Using configurations of left and right circularly polarized light, the graviton chirality is measured to be negative for the $n/(2n{+}1)$ states while it is positive for the $n/(2n{-}1)$ states [37]. Within the parton description, the chirality of the graviton [74] is negative (positive) if the effective magnetic field seen by the parton hosting the exciton is parallel (antiparallel) to the applied field [34].
Discussion. The CFEs should in principle be observable by light scattering. We expect that for each CFE mode, $S_{\mbox{\boldmath$q$}}$ is maximized at wave vectors where its energy intersects the GMP mode. Thus, in place of a single GMP mode, we expect many discrete modes, highlighted in the schematic Fig. 1. Evidence of multiple modes has been seen in numerical studies [75, 24]. For a disorder-free system Raman scattering probes only small $q\ell$ modes, but an external one-dimensional periodic potential, e.g., that produced by a surface grating or by piezoelectric coupling to a surface phonon, can help select modes with finite wave vectors [15, 16] and should reveal the discrete CF modes. Larger $q\ell$ can also be accessed in tilted field Raman experiments at low electron densities. While there is only one graviton for $\nu{=}1/3$, its splitting into two at small but non-zero $q$ was observed in Ref. [13], and explained in Ref. [24] in terms of CFEs.
@@
Acknowledgments. We acknowledge useful discussions with Abhishek Anand, Rakesh Dora, Lingjie Du, Mytraya Gattu, Zlatko Papić, Songyang Pu, and Arkadiusź Wójs. This work was undertaken on the Nandadevi supercomputer maintained and supported by the Institute of Mathematical Science’s High-Performance Computing Center, India. Some of the numerical calculations were performed using the DiagHam libraries [78]. ACB thanks the Science and Engineering Research Board (SERB) of the Department of Science and Technology (DST) for funding support via the Mathematical Research Impact Centric Support (MATRICS) Grant No. MTR/2023/000002. JKJ was supported in part by the U. S. Department of Energy, Office of Basic Energy Sciences, under Grant No. DE-SC-0005042. We also thank National Supercomputing Mission (NSM) for providing computing resources of ‘PARAM Brahma’ at IISER Pune, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and Department of Science and Technology (DST), Government of India.
@@
This supplemental material contains certain technical details that are not included in the main text. In Sec. SI we present the method of and results obtained from direct projection. Energies of the GMP density wave mode obtained from acting the projected density operator on the exact LLL Coulomb ground state are presented in Sec. SII. (In Fig. 3 of the main text we presented GMP energies obtained from acting the projected density operator on the model Laughlin/Jain ground states).
SI Projection into LLL: Method and results

### v2

@@ Abstract
Introduction. A dramatic manifestation of strong correlations is “fractionalization of the electron,” which refers to the notion that the emergent elementary particles (that is, the weakly interacting degrees of freedom) carry quantum numbers of a split electron [1]. For example, when an electron is added to a Luttinger liquid, its spin and charge separate [2], and when a fully spin-polarized electron is added to an incompressible FQH state, it disintegrates into fractionally-charged excited CFs [3, 4]. The emergent particles have their own independent dynamics and provide a natural explanation of the otherwise puzzling phenomenology.
This Letter concerns the Girvin-MacDonald-Platzman (GMP) density wave mode of the FQH states [5, 6]. Since the work of GMP numerous experimental [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 16] and theoretical [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39] studies have addressed the neutral excitations of the FQH states. Very recent inelastic light scattering experiments by Liang et al. [40] with circularly polarized photons have measured the chirality of the long wavelength neutral excitation, which has been called “graviton” because of its spin-2 quadrupolar character in a geometric interpretation, and found it to be consistent with theoretical predictions [26, 32, 41].
@@
$S(\mbox{\boldmath$q$},E)=\sum_{\alpha}|\langle\Psi_{\nu}^{\alpha}|\Psi^{\rm GMP}_{\mbox{\boldmath$q$}}\rangle|^{2}\delta(E_{\alpha}-E),$ (1)
@@
Here $S_{\mbox{\boldmath$q$}}$ is the LLL-projected static structure factor. In general, $E^{\rm SMA}_{\mbox{\boldmath$q$}}$ is the “average energy” of the excitations coupled to the ground state by $\bar{\rho}_{{}_{\mbox{\boldmath$q$}}}$. $\Psi^{\rm GMP}_{\mbox{\boldmath$q$}}$ can be evaluated explicitly for small systems for which the Fock space representation of the exact or the trial ground state is available but not for larger systems. GMP related the energy of this mode to the unprojected static structure factor, which is the Fourier transform of the pair correlation function $g(\mbox{\boldmath$r$})$, which can be evaluated for a trial wave function for fairly large systems [42, 43, 44, 45], allowing for a determination of the energy of the GMP mode [5, 6, 46, 22].
@@
We employ two methods for treating the LLL projection. In “direct projection” one expresses the wave function in the Fock basis and retains only the part that resides in the LLL [18]. This method can be implemented only for small systems, as the dimension of the Fock space grows exponentially with $N$ [see Supplemental Material (SM) [50]]. The Jain-Kamilla (JK) method [51, 52, 53, 54, 55] allows treatment of much larger systems. The two methods generally produce wave functions that are extremely close but not identical.
Figure 1: (a) GMP density wave mode, created by applying LLL-projected electron density operator on the ground state. (b) Jain CF ground state at $\nu{=}2/5$. (c)-(g) Examples of single CF excitons. (h),(i) Schematic depiction of splitting of the GMP density wave mode (dashed black curve) into a ladder of CF excitons (blue dotted) which can be obtained by diagonalizing the Coulomb interaction in the CFE basis. The actual lowest neutral mode is well described as the primary CFE shown in (c). For $\nu{=}n/(2n{\pm}1)$ the GMP graviton is essentially identical to the primary CF graviton (h) while for $\nu{=}n/(4n{\pm}1)$ with $n{>}1$ the GMP graviton splits into two CF gravitons, which can be also be nicely understood as the primary CF and parton gravitons (i). (j) Parton construction of $\nu{=}2/9$. Each electron is divided into one parton of charge $e/9$ filling two LLs and four partons of charge $92e/$ 9, each filling a single LL. (k) CF exciton. (l) Parton exciton.
Spherical geometry. We use the spherical geometry [56] in which $N$ electrons move on the surface of a sphere in the presence of a radial magnetic field generated by a magnetic monopole producing a total flux $2Qhc/e$, where $2Q$ is an integer. The LL orbitals are given by the monopole harmonics $Y_{Q,l,l_{z}}$ [57, 58], where $l{=}|Q|,|Q|{+}1,{\cdots}$ and $l_{z}{=}{-}l,{-}l{+}1,{\cdots},l$. The $n$th LL corresponds to $l{=}|Q|{+}n$, with $n{=}0,1,2{\cdots}$. The eigenstates are labeled by the total orbital angular momentum $L$ and its $z$-component $L_{z}$.
@@
In all cases studied, the GMP mode and CFEs constructed with direct or JK projection vanish identically for $L{=}1$. For higher $L$, we have studied the following ($\nu$, $N$) systems with direct projection: ($1/3$,6) for $L{\leq}7$; ($1/3$,8) for $L{\leq}7$; ($2/5$,6) for $L{\leq}4$; ($2/5$,8) for $L{\leq}11$; ($2/3$,6) for $L{\leq}4$; ($2/3$,8) for $L{\leq}4$; ($2/7$,6) for $L{\leq}4$; ($2/7$,8) for $L{\leq}4$; ($2/9$,6) for $L{\leq}4$; ($2/9$,8) for $L{=}2$; bosonic ($2/3$,6) for $L{\leq}4$; bosonic ($2/3$,8) for $L{\leq}4$. We find ${\cal O}{=}1$ in arbitrary precision calculations for all of these. This is nontrivial, because, as shown in Table 1, the number of CFEs at a given $L$ is $N$-independent (provided $N$ is sufficiently large) and much smaller than the number of basis functions, which ranges to several thousand for the systems studied and increases exponentially with $N$. This gives us confidence that the GMP mode is generally contained entirely within the (direct-projected) CFE subspace, although we have not yet succeeded in devising a general proof.
@@
The values of ${\cal O}_{\rm JK}$ obtained for the JK-projected CFEs are given in Table 1 for $\nu{=}n/(2n{+}1)$. In all cases studied, we find that ${\cal O}_{\rm JK}{>}0.997$, demonstrating that Eq. (4) is practically exact for the JK projection. The JK-projected CFEs for $n/(4n{\pm}1)$ Jain states are considered below.
The relation in Eq. (4) between the GMP mode and single CFEs is unexpected as these are motivated by qualitatively different physical descriptions. The GMP ansatz creates a coherent superposition of states containing a single electron-hole pair while a CFE involves the excitation consisting of a single CF particle and a single CF hole, each of which, in isolation, has a fractional charge of magnitude $e/(2n{\pm}1)$ [3]. Indeed, the dispersion of the GMP mode is generally rather different from that of the primary CFE [48, 49, 22, 62]. Moreover, because an electron added to a FQH state is described as a complex bound state of $12n{\pm}$ 1 excited CFs [63, 64, 65, 66], one would a priori expect an electron-hole pair of the GMP mode to involve $12n{\pm}$ 1 CFEs. Eq. (4) is far from obvious also from energetic considerations (see SM [50]).
CFE ansatz for $S(\mbox{\boldmath$q$},E)$. Substituting Eq. (4) into Eq. (1) suggests that the SMA of Eq. (2) can be replaced by a “CFE Ansatz” for the LLL-projected DSF:
@@
Figure 2: The energies of the GMP mode (green diamonds), the JK-projected primary CF exciton (blue dots) for $\nu{=}1/3,~{}2/5,~{}3/7,~{}1/5,~{}2/9$ and $2/7$, and the JK-projected primary parton exciton (black pentagrams) for $\nu{=}2/7,2/9$. Red dashes show the exact Coulomb spectrum. The spherical geometry is used, and the GMP energy is computed using the Laughlin (Jain) ground state for $1/3,~{}1/5$ ($2/5,~{}3/7,~{}2/9,~{}2/7$).
Figure 3: Plot of $(\rho(r){-}\rho_{0})/\rho_{0}$ for the JK-projected CF graviton with $L{=}2{=}L_{z}$ at $\nu{=}1/3$, $2/5$ and $3/7$, where $\rho(r)$ and $\rho_{0}$ are the densities of the graviton and the ground state, and $r$ is the distance from the center.
@@
For the Jain states at $\nu{=}n/(4n{\pm}1)$ with $n{>}1$ the energy of the GMP graviton differs significantly from that of the primary CF graviton, as seen in Fig. 2. Fig. 4 shows the spectral weight obtained by diagonalizing the Coulomb interaction within the direct-projected CF graviton space (which has three linearly independent modes). It is in excellent agreement with that obtained by ED, supporting the validity of Eq. (5). The presence of two prominent peaks indicates that the GMP graviton splits into two CF gravitons for a range of broadening.
Remarkably, Nguyen and Son [71] predicted an additional high-energy graviton at $n/(2pn{\pm}1)$ for $22p{\geq}4,n{\geq}$ 2, motivated by the observation that a Dirac CF field theoretical description of $\nu{=}n/(4n{\pm}1)$ states by Goldman and Fradkin [72] and Wang [73] does not satisfy a bound derived by Haldane [74] which says that the coefficient of the $q^{4}$ term in the projected static structure factor is lower-bounded by $({\cal S}{-}1)/8$, where ${\cal S}$ is the Wen-Zee shift [75]. Balram et. al. [37] proposed to identify the second, higher energy graviton as a “parton graviton.” In the parton construction of the FQH states  [76] one divides each electron into an odd number of unphysical fractionally charged fermions called partons [see Fig. 1(j)], places each species of partons in an integer quantum Hall state, and then combines the partons back into physical electrons. In the Jain $n/(2pn{\pm}1)$ states, one species of partons forms the $\nu{=}{\pm}n$ state and remaining $2p$ partons the $\nu{=}1$ state. Balram et. al. [37] defined the parton exciton as the exciton in $\Phi_{1}$ i.e.,
@@
Surprisingly, Eq. (4) is not a good approximation for the JK projected CFEs for the Jain $\nu{=}n/(4n{\pm}1)$ states with $n{>}1$. For example, for $\nu{=}2/7$ ($\nu{=}2/9$) we find ${\cal O}_{\rm JK}{=}0.5076$ (${\cal O}_{\rm JK}{=}0.9191$) for $N{=}8$ at $L{=}2$. [To project, we write $\Psi_{n/(4n{\pm}1)}{=}\Phi_{1}^{2}\Psi_{n/(2n{\pm}1)}$; 2/7 has a single CF graviton because 2/3 does (the hole conjugate of the single graviton at 1/3), and 2/9 inherits 2 CF gravitons from 2/5.] It turns out that the GMP graviton lies, to an excellent approximation, within the 2-dimensional subspace defined by the primary CF and parton gravitons. For the $N{=}8$ particle $\nu{=}2/7$ ($\nu{=}2/9$) state, the GMP graviton has ${\cal O}_{\rm JK}{=}0.9995$ (${\cal O}_{\rm JK}{=}0.9986$) with this subspace. Using configurations of left and right circularly polarized light, the graviton chirality is measured to be negative for the $n/(2n{+}1)$ states while it is positive for the $n/(2n{-}1)$ states [40]. Within the parton description, the chirality of the graviton [78] is negative (positive) if the effective magnetic field seen by the parton hosting the exciton is parallel (antiparallel) to the applied field [37].
Discussion. The CFEs should in principle be observable by light scattering. We expect that for the $\alpha$ CFE mode, $S^{\rm CFE}_{\mbox{\boldmath$q$},\alpha}$ is maximized at wave vectors where its energy intersects the GMP mode. Thus, in place of a single GMP mode, we expect many discrete modes, highlighted in the schematic Fig. 1. Evidence of multiple modes has been seen in numerical studies [79, 25]. For a disorder-free system Raman scattering probes only small $q\ell$ modes, but an external one-dimensional periodic potential, e.g., that produced by a surface grating or by piezoelectric coupling to a surface phonon, can help select modes with finite wave vectors [15, 16] and should reveal the discrete CF modes. Larger $q\ell$ can also be accessed in tilted field Raman experiments at low electron densities. While there is only one graviton for $\nu{=}1/3$, its splitting into two at small but non-zero $q$ was observed in Ref. [13] and explained in Ref. [25] in terms of CFEs.
@@
Acknowledgments. We acknowledge useful discussions with Abhishek Anand, Rakesh Dora, Lingjie Du, Mytraya Gattu, Zlatko Papić, Songyang Pu, and Arkadiusź Wójs. This work was undertaken on the Nandadevi supercomputer maintained and supported by the Institute of Mathematical Science’s High-Performance Computing Center, India. Some of the numerical calculations were performed using the DiagHam libraries [82]. ACB thanks the Science and Engineering Research Board (SERB) of the Department of Science and Technology (DST) for funding support via the Mathematical Research Impact Centric Support (MATRICS) Grant No. MTR/2023/000002. JKJ was supported in part by the U. S. Department of Energy, Office of Basic Energy Sciences, under Grant No. DE-SC-0005042. We thank the National Supercomputing Mission (NSM) for providing computing resources of ‘PARAM Brahma’ at IISER Pune, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and the Department of Science and Technology (DST), Government of India.
@@
This supplemental material contains certain technical details that are not included in the main text. In Sec. SI we provide a primer on the Girvin-MacDonald-Platzmann (GMP) and composite-fermion exciton (CFE) modes. Then, in Sec. SII we present the method of and results obtained from direct projection. Energies of the GMP density wave mode obtained from acting the projected density operator on the exact LLL Coulomb ground state are presented in Sec. SIII. (In Fig. 3 of the main text we presented GMP energies obtained from acting the projected density operator on the model Laughlin/Jain ground states.). Sec. SIV displays the effect of disorder-broadening or the resolution of the measurement on the spectral function. The last section, Sec. SV, presents the dynamical structure factor for total orbital angular momentum $L{\geq}2$ at $\nu{=}2/5$.
SI Background
Density operator
The density operator in the first quantized notation is given by
$\rho(\theta,\phi)=\sum_{i=0}^{N}\delta(\cos\theta-\cos\theta_{i})\delta(\phi-\phi_{i})=\sum_{l=0}^{\infty}\sum_{m=-l}^{l}Y^{*}_{0lm}(\theta,\phi)\sum_{i=1}^{N}Y_{0lm}(\theta_{i}\phi_{i})$ (S1)
where $Y_{0lm}$, a special case of the “monopole harmonics” $Y_{qlm}$, are the same as the usual spherical harmonics $Y_{l,m}$. Identifying this as a mode expansion in spherical harmonics we define the $l,m$ mode of the density operator as
$\rho_{L=l,L_{z}=m}=\sum_{i=1}^{N}Y_{0lm}(\theta_{i}\phi_{i})$ (S2)
It is easy to check that $\rho_{lm}$ adds an angular momentum $L{=}l,~{}L_{z}{=}m$ to the state it acts on. In particular, it produces a many-body state with $L{=}l$ and $L_{z}{=}m$ when it acts on the ground state (which is a rotation-invariant state $L{=}0$).
Since $Y_{000}{\sim}1$, action of $\rho_{0,L{=}0,m{=}0}$ on the ground state returns the ground state. In general, the spectral functions being calculated here are rotation invariant, therefore it is sufficient for our purposes to consider the lowest weight state ($L_{z}{=}{-}L$) of each angular momentum multiplet. When the spherical harmonics are expressed in terms of the spinor coordinates $u_{i}{=}\cos(\theta_{i}/2)e^{i\phi/2},~{}v_{i}{=}\sin(\theta_{i}/2)e^{{-}i\phi_{i}/2}$, the highest weight density operators have the form $Y_{Q,L,L_{z}}(u_{i},v_{i}){\propto}v^{L}_{i}\bar{u}^{L}_{i}$. The corresponding LLL projected density operator is obtained by replacing $\bar{u}$ with $\partial/\partial_{u}$ and normal-ordering the derivatives, i.e., moving all the derivatives to the left. With the explicit form of the projected density operator, one can then show that the $L{=}1$ GMP mode is annihilated upon projection to the LLL [62, 59]. Thus the GMP mode starts from $L{=}2$ (graviton) and extends up to $L^{\rm GMP}_{\rm max}{=}2Q{\sim}N/\nu$ for large $N$ because the GMP operator creates a coherent superposition of states containing an electron and a hole in the LLL each of which carries a single-particle angular momentum of $l{=}Q$.
For the CFE state defined in Eq. (3) of the main text, $\Phi_{n}$, the wave function of $n$ filled LLs is constructed at an effective flux of $2Q^{*}{=}2Q{-}2(N{-}1){=}N/n{-}n$. There are many CFE modes corresponding to excitations in $\Phi_{n}$ across different LLs. The lowest energy CFE called the “primary CFE”, corresponds to the $n{-}1{\rightarrow}n$ excitation. The angular momenta of the constituent CF particle and CF hole are $l^{\rm CFP}{=}Q^{*}{+}n$ and $l^{\rm CFH}{=}Q^{*}{+}n{-}1$. The primary CFE extends from $L^{\rm CFE}_{\rm min}{=}2$ (the $L{=}1$ CFE is also annihilated upon projection into the LLL [18]) to $L^{\rm CFE}_{\rm max}{=}l^{\rm CFH}{+}l^{\rm CFP}{=}2Q^{*}{+}2n{-}1{=}(N/n{+}n{-}1){\sim}N/n{=}N/\nu^{*}$ for large $N$, where $\nu^{*}$ is the effective filling of CFs [34]. In general, the $m{\rightarrow}m{+}K$ CFE mode starts from $L{=}K$ (for certain CFEs the $L{=}K$ excitation is annihilated upon projection to the LLL [18, 68] in which case the mode starts from $L{=}K{+}1$) and extends up to $2Q^{*}{+}2m{+}K$ since the CF particle and CF hole angular momenta are $l^{\rm CFP}{=}Q^{*}{+}m{+}K$ and $l^{\rm CFH}{=}Q^{*}{+}m$. As an example, at $L{=}2$, there are at most two CFEs for $\nu{=}1/3$ [$0{\rightarrow}1$ and $0{\rightarrow}2$ excitons (Note that the $0{\rightarrow}3$ exciton has $L{\geq}3$.)], and at most three for $\nu{=}2/5$ ($1{\rightarrow}2$, $1{\rightarrow}3$ and $0{\rightarrow}2$). An explicit construction of the wave functions is required to get the final counting of states at a given $L$ since certain CFEs can be annihilated upon projection to the LLL. The states at $\nu{=}(n{+}1)/(2n{+}1)$ are related by particle-hole symmetry to the states at $\nu{=}n/(2n{+}1)$ [83]. Thus, the spectra at the two fillings, particularly their gravitons, can be related to each other. For example, the 2/3 state has only one graviton since its hole conjugate at 1/3 supports only a single graviton. Furthermore, the chirality of the graviton at $(n{+}1)/(2n{+}1)$ is opposite to the chirality of the graviton at $n/(2n{+}1)$ [78, 37, 38]. Since the secondary Jain states at $n/(4n{\pm}1)$ can be viewed as a product of the bosonic $1/2$ Laughlin and $n/(2n{\pm}1)$ Jain states, the $n/(4n{\pm}1)$ state inherits (by inherit, here we mean if a LLL FQH ground state wave function $\Psi$ can be written as a product of two LLL FQH ground state wave functions $\Psi^{\alpha}$ and $\Psi^{\beta}$, and $\Psi^{\alpha}$ and $\Psi^{\beta}$ have some particle-hole like modes, we generally expect $\Psi$ to inherit them) the gravitons of the $n/(2n{\pm}1)$ Jain state (additional graviton inherited from the bosonic $1/2$ Laughlin state can also be present).
Before proceeding to the results of our calculations, we ask whether the relation given in Eq. (4) of the main text may be anticipated from energetic considerations. Without explicit calculation, the extent to which the GMP mode and the CF excitons (CFEs) describe the low-energy excitations is unknown. Calculations show that the GMP mode is close to the lowest energy mode of the $n/(2n{\pm}1)$ states at small wave vectors but not at large wave vectors, while for the $n/(4n{\pm}1)$ states it is not a low energy mode at any wave vector (except for the $1/5$ state). The lowest energy CFE has been found to provide an excellent description of the lowest energy mode for all of these fractions at all wave vectors. The higher energy excitations are described by multiple CFEs; for example, excitations with 3 CF cyclotron energy can be produced by creating a single CFE, two CFEs, or 3 CFEs. From this perspective, it is not obvious, a priori, that the GMP mode can be expressed entirely in terms of single CFE basis states.
SII Direct projection into LLL: Method and results